OpenStudy (anonymous):

Suppose that a balanced die is rolled repeatedly until the same number appears on two successive rolls, and let X denote the number of rolls that are required. Determine the value of Pr(X=x) for x=2,3,...

OpenStudy (jamesj):

Well, what is Pr(X=2)? And Pr(X=3)?

myininaya (myininaya):

what is a balanced die that is a regular die right?

OpenStudy (anonymous):

Its just a die that isn't loaded...and doesnt favor one number over the other...

OpenStudy (jamesj):

Pr(X=2) = Pr(second roll = first roll) = 1/6 Now Pr(X=3) = Pr(2nd roll not = first roll).Pr(3rd roll=2nd roll) = 5/6 . 1/6 Hence Pr(X=4) = ...

OpenStudy (jamesj):

I like this question a lot because you'll see this: Once you find the answer, you'll also be able to show that \[\sum_{n} \Pr(X=n) \ = 1\] which makes sense as the probability that infinity many rolls have no repeats is zero.

OpenStudy (anonymous):

My friend got the answer: (5/6)^(x-2) times (1/6)

OpenStudy (anonymous):

I came to that solution I believe...but it's been taking me a long time to realize how I got it...

OpenStudy (jamesj):

who cares. What's the logic for the answer? Follow what I wrote above and hopefully the first two cases make sense. What is Pr(X=4)?

OpenStudy (anonymous):

(5/6)^2 (1/6) ?

OpenStudy (jamesj):

Pr(X=4) = Pr(2nd roll not = 1st roll) . Pr(3rd roll not = 2nd roll) . Pr(4th roll = 3rd roll) = 5/6 . 5/6 . 1/6

OpenStudy (jamesj):

right. And now generalize

OpenStudy (anonymous):

Oh thank you so very much! I understand the problem quite well now...or I hope I do. Are you a math major by the way?

OpenStudy (jamesj):

was, yes

OpenStudy (anonymous):

That was a very helpful way of showing me how to understand the problem, thank you again.